A jogger starts at position A and runs 70m due west to position B. He then turns around and runs 85m due east to position C Define his final displacement when using: Position A as a reference point Position B as a reference point Determine the total distance travelled by the jogger.
1
Expert's answer
2021-10-25T15:36:48-0400

A particle’s position is the location of the particle with respect to a chosen 

reference point that we can consider to be the origin of a coordinate system. You can freely choose the positive and negative direction. The displacement of an object is defined as it’s change in position in some 

time interval: 

x = 70 - 85 = - 15 m

Total distance = 70 + 85 = 155 m